y=-1
-2x-2y=-2 solve by substitution

Answer by waynest(281)

y = -1
-2x - 2y = -2
substitute -1 for y in the second equation:
-2x - 2(-1) = -2
-2x + 2 = -2
-2x + 2 - 2 = -2 - 2
-2x = -4
-2x/-2 = -4/-2
x = 2
substitute 2 for x in the second equation:
-2(2) - 2y = -2
-4 - 2y = -2
-4 + 4 - 2y = -2 + 4
-2y = 2
-2y/-2 = 2/-2
y = -1